OpenStudy (anonymous):

find the antiderivative of F''=24x^2+2x+10 F(4)=5, F'(1)=-3

OpenStudy (dumbcow):

use power rule x^n --> x^(n+1)/(n+1) F' = 8x^3 + x^2 +10x + C F'(1) = -3 8+1+10+C = -3 C = -22 F = 2x^4 +x^3/3 +5x^2 -22x +C F(4)=5 2*256 + 64/3 + 80 - 88+C = 5 C = -1561/3 F = 2x^4 + x^3/3 +5x^2 -22x - 1561/3

OpenStudy (anonymous):

dumbcow please come look at my latest question when you are done, thanks you helped alot